If the larger number is subtracted from 6 times the smaller number, the result is 20. If twice the larger number is added to 4 times the smaller, the result is 56. Find the two numbers
Answer by checkley77(12844)

6Y-X=20 MULTIPLY THIS EQUATION BY 2 & ADD.
4Y+2X=56
12Y-2X=40
-------------
16Y=96
Y=96/16
Y=6 ANS.
6*6-X=20
36-X=20
-X=20-36
-X=-16
X=16 ANS.
PROOF:
4*6+2*16=56
24+32=56
56=56